Coronado Company acquired a plant asset at the beginning of Year 1. The asset has an estimated service life of 5 years. An employee has prepared depreciation schedules for this asset using three different methods to compare the results of using one method with the results of using other methods. You are to assume that the following schedules have been correctly prepared for this asset using (1) the straight-line method, (2) the sum-of-the-years-digits method, and (3) the double-declining-balance method. Sum-of-the Double-Declining- Year Straight-Line Years-Digits Balance $9,360 9,360 9,360 9,360 9,360 $46,800 $15,600 12,480 9,360 6,240 3,120 546,800 $20,800 12,480 7,488 4,493 1,539 $46,800 Total Answer the following questions. Your answer is correct. What is the cost of the asset being depreciated? jszo00 Cost of asset Attempts: 1 of 3 used What amount, if any, was used in the depreciation calculations for the salvage value for this asset? Salvage value Attempts: 0 of 3 used
Answer #1

B. Cost of asset = $ 52000

Total Value depreciated in all three methods = $ 46,800

So Salvage Value = 52000-46800 = $ 5200

C . As double declining method uses the double of a rate of straight-line method.

and we can also see from the table that in year 1 there is the highest depreciation in double declining balance method.

So answer is Double Declining method

D. We can see from the table that in Year 4 there is the highest depreciation in Straight line method. Thus the method that produces the highest charge to income in Year 4 will be the Straight Line Method.

E. We will calculate 3rd year-end book value of the asset under all three method.

1. Straight line = 52000-9360-9360-9360 = 23,920

2. Sum of Year's digit method = 52000-15600-12480-9360 = 14,560

3. Double declining balance method = 52000-20800-12480-7488 = 11,232

So we can see that in Straight line method produces highest book value for the asset at the end of the year 3.

F. From point E we have seen that under Double-declining balance method, value of the asset at end of the Year 3 is lowest, So if we will sale it, we will gain more.

So Under Double decling balance method we will gain maximum profit.